OpenStudy (anonymous):

Find a quadratic function to model the values in the table. Predict the value of y for x = 6. X Y 1---2 0---2 3---(-10) 6---?

OpenStudy (anonymous):

I know the y-intercept is 2 because of the point (0,2)

OpenStudy (anonymous):

\[y=ax^2+bx+c\] a+b+c=2 c=2 9a+3b+c=-10 a+b=0--->3a+3b=0 9a+3b=-10-c=-12 6a=-12---->a=-2 b=-a=2 \[y=-2x^2+2x+2\] y(x=6)=-2·36+2·6+2=-72+12+2=-58

OpenStudy (anonymous):

Ok Im confused. How did u get all that with the 9a + 3b

OpenStudy (anonymous):

i get it now after staring at it

OpenStudy (anonymous):

First three lines are the equation system you have to solve a+b+c=2 c=2 9a+3b+c=-10